Solve by completing the square and then tell which form is x^(2)-10x+11
form is either (x-?)^(2)=? or
(x+?)^(2)=?

Answer by oscargut(2103)

x^(2)-10x+11 =0
x^2-10x+25 = 14
(x-5)^2= 14
Form is: (x-5)^(2)=14
Solutions: sqrt(14)+5 and -sqrt(14)+5